In a class test containing 10 questions, 5 marks are awarded for every correct answer, and (–2) marks are awarded for every incorrect answer and 0 for questions not attempted.
 Heena gets two correct and five incorrect answers out of seven questions she attempts. What is her score?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question,
Marks awarded for 1 correct answer = 5
Then,
Total marks awarded for 2 correct answer = 2 × 5 = 10
Marks awarded for 1 wrong answer = -2
Then,
Total marks awarded for 5 wrong answer = 5 × -2 = -10
Marks awarded for questions not attempted is = 0
∴Total score obtained by Heena = 10 + (-10)
= 10–10
=0

Thanks